Construct the confidence interval for the population variance for the given values. Round your answers to one decimal place. n=27, s^2 =10.5 and c=0.9
Answer #1

Answer:

C= 90% = 0.90,  =0.1

n=27, df =n-1 = 27-1 = 26

formula for confidence interval is

now calculate both critical values with df = 26

using X2 table we get critical value as

X20..05=38.8851

X20. 95 =15.3792  

= (7.0207 < < 17.7513)

Therefore, based on the data provided, the 90% confidence interval

for the population variance is= 7.0<   <17.8

confidence interval is = ( 7.0 , 17.8 )